OpenStudy (anonymous):

What is the quotient of 3 1/3 divided by 5/7 in simplest form? three and five over twenty-one three and seven over fifteen four and two over three four and eight over twenty-one

OpenStudy (anonymous):

@happinessbreaksbones

OpenStudy (anonymous):

@JaziBoo

OpenStudy (anonymous):

and u write out the numbers, the way it is written is kinda confusing :P

OpenStudy (anonymous):

*can

OpenStudy (anonymous):

I got it fixed at the top in number form

OpenStudy (anonymous):

Otay, imma check it out now :)

OpenStudy (anonymous):

\[3 \frac{ 1}{ 3} \frac{ 5 }{ 7 }\]

OpenStudy (anonymous):

I'd say its the third one :) 4 2/3

OpenStudy (anonymous):

\[\frac{ 5 }{ 8 }\times20\]

OpenStudy (anonymous):

wat?? :P

OpenStudy (anonymous):

Im asking another question LOL

OpenStudy (anonymous):

its definately the third choice, i worked it out ;) :P

OpenStudy (anonymous):

oh ok lmaoo :P

OpenStudy (anonymous):

i think the second one u asked is 25 over 2 idk :P

OpenStudy (anonymous):

Thx

OpenStudy (anonymous):

(_8|)

OpenStudy (anonymous):

u welcumzzz :D

OpenStudy (tkhunny):

Very clean! \(\dfrac{3\;1/3}{5/7} = \dfrac{10/3}{5/7} = \dfrac{10}{3}\cdot\dfrac{7}{5} = \dfrac{10}{5}\cdot\dfrac{7}{3} = 2\cdot\dfrac{7}{3} = \dfrac{14}{3} = 4\;\dfrac{2}{3}\) Don't guess. Prove!